OpenStudy (anonymous):

Please Help!!! :( Mr. Grant made a triangular playhouse for his children in one corner of the family room. The triangular base has a height of 1.6 m and a base of 2.1 m. How many square meters of plywood did Mr. Grant use if the playhouse is 1.8m tall and 2.6m wide.

Directrix (directrix):

Do you have a diagram of the playhouse? Is it a triangular prism?

OpenStudy (anonymous):

no this is the problem

Directrix (directrix):

I can't say for sure that this is how the playhouse looks but it is my best guess.

OpenStudy (anonymous):

okay thanks but how do i solve it?

Directrix (directrix):

Here's my best shot on the problem from the given information. The wood needed would be a measure of surface area. Playhouse wood = 3 rectangular walls + 2 bases (floor and roof) P.W. = 3 * 1.8 * 2.6 + 2* (1/2)* (2.1)*(1.6) Plywood Needed = 17.4 square meters

OpenStudy (anonymous):

okay thank you very much!

Directrix (directrix):

Well, I can't be sure it is correct because I don't know how the figure looks.

OpenStudy (anonymous):

its okay. can you help me with this problem: A garden 40 meters wide and 72 meters long is to be enclosed by a fence. At $8.90 per meter, what will the fencing cost?

Directrix (directrix):

We'll divide up the problem. Do you want to find the perimeter of the garden or the cost of the fence after I find the perimeter of the garden? Your choice. :)

OpenStudy (anonymous):

i'll find the perimeter

OpenStudy (anonymous):

is the perimeter 224m?

Directrix (directrix):

2 * 72 + 2 * 40 = 224 --> perimeter so 224 * $8.90 = $ 1,993.60

OpenStudy (anonymous):

okay thank you :). and this will be the last question. Part of the edging around a regular decagon having a perimeter of 213 cm was used to edge a regular octagon. If the sides of the octagon are the same length as those of the decagon, how much edging was used for the octagon?

Directrix (directrix):

You figure out the how long each side of the decagon edging is. Then, I'll do the octagon part. Or, we could flip. Your choice.

OpenStudy (anonymous):

I don't know how to figure out the lengths

Directrix (directrix):

How many sides does a decagon have?

OpenStudy (anonymous):

10

Directrix (directrix):

If 213 feet were used to edge (same as getting its perimeter) and the decagon's sides are all the same (that's what regular means in the problem), then what would you do to the perimeter to get the length of one side of the decagon?

OpenStudy (anonymous):

213(4)?

Directrix (directrix):

213 divided by 10 = 21.3 cm per each of the ten sides. An octagon has 8 sides and if all are the same length, then 8 times 21.3 cm was used to edge the octagon flower bed. 21.3 * 8 = 170.4 cm. You could just take 8/10 of the perimeter of the decagon and get the same result: .8 * 213 = 170.4 cm.

OpenStudy (anonymous):

okay thank you